OpenStudy (anonymous):

In a lottery, there are 2 000 000 tickets sold. The prizes are as follows: $1 000 000/ 1prize , $50 000/5 prizes, $1000/ 10 prizes, and $50/ 50 prizes. What should the lottery operators charge per ticket to make a 40% profit?

OpenStudy (anonymous):

The answer is 1.05 $

OpenStudy (anonymous):

with this i dont get the answer to be 1.05

OpenStudy (anonymous):

(Money In - Money Paid Out) / Money In = 0.4 Money in is 2 000 000 x where x is the charge per ticket Money Paid Out = 1 000 000 + 5*50 000 + 10*1000 + 50*50 = 1262500 (2 000 000 x - 1 262 500) / 2 000 000 x = 0.4 Solve for x.

OpenStudy (anonymous):

dont the x's cancel out when u divide?